OpenStudy (anonymous):

Please help! Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______ Graph of f of x and g of x. f of x equals 1 over 3 x minus 2 and g of x equals 1 over 3 x plus 3. Numerical Answers Expected!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

That was my reaction .-.

OpenStudy (aum):

Are these the given functions? \(\large f(x) = \frac{1}{3x}-2\) \(\large g(x) = \frac{1}{3x}+3\)

OpenStudy (anonymous):

I believe so yeah

OpenStudy (aum):

\[ \large g(x) = f(x) + k \\ \large k = g(x) - f(x) = \frac{1}{3x}+3 - (\frac{1}{3x}-2) = ? \]

OpenStudy (anonymous):

Okay..